A projectile is dropped from a tower. Distance travelled by the projectile in the last two seconds is 88m. What is the height of the tower?

Grade:12

3 Answers

Soumya
6 years ago
Well distance travelled by projectile in last 2 second =88mSo from this we can find out the velocity at that point that is 44m/s and its vertical component is also 44m/sUsing 3rd law of motion we ge h=44*44/2*10=96.8
Sushant Singh Raghu
6 years ago
Calculate this in two steps.1) 1st for the journey H-80 with tine t-2 and then for the last 2 seconds.
 
 
sam
6 years ago
ANS. 151.25
2 equations
H=1/2 g t^2
H-88= ½ g (t-2)^2
2 equations 2 unknowns, solve it, find H
on solving, we get H=151.25 m